Find the wavelength of the third line in the Lyman series, and identify the type of EM radiation.

I need to know what the governing principles are for this equation.
Answer #1

Wave number of spectral lines of hydrogen spectrum, 1/λ = R {1/n12 – 1/n22}

Rydberg constant, R = 1.097x 107 m-1

For Lyman series, n1 = 1 and n2= 2, 3, 4, 5, ………

For the third line: n1 = 1 and n2 = 4

1/λ = 1.097 x 107 {1/12 – 1/42} = 1.028 x 107

Wavelength of third line of Lyman series, λ = 972.7 x 10-10 m

Radiation of this wavelength belongs to ultraviolet region of EM radiation.

The equation for the wave number is derived using Bohr’s postulates of hydrogen atom. According to Bohr’s postulates, an atom emits radiation whenever an electron jumps from a higher orbit to a lower orbit.